OpenStudy (anonymous):

7^2x-1= 316. exponential equations...

OpenStudy (anonymous):

\[b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (anonymous):

\[7^{2x-1}=316\iff 2x-1=\frac{\ln(316)}{\ln(7)}\]

OpenStudy (anonymous):

use a calculator to compute the right hand side, then solve for \(x\) by adding one and dividing the result by 2

OpenStudy (anonymous):

let me know what you get as a final answer

OpenStudy (anonymous):

i got 1.48

OpenStudy (anonymous):

then there is a mistake somewhere

OpenStudy (anonymous):

i did what you said to do

OpenStudy (anonymous):

here is \[\frac{\ln(316)}{\ln(7)}\] http://www.wolframalpha.com/input/?i=ln%28316%29%2Fln%287%29 is the what you got ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

about 2.95866 add 1 to get 3.95866 then divide by 2

OpenStudy (anonymous):

perhaps you put it in the calculator wrong you have to close the parentheses after the log on top

OpenStudy (anonymous):

i got 1.979

OpenStudy (anonymous):

that is it